OpenStudy (anonymous):

Assuming the balls to be identical except for difference in colours, the number of ways in which one or more balls can be selected from 10 white, 9 green and 7 black balls is :

OpenStudy (dumbcow):

not sure of a general formula here...but there is a pattern your question is vague, so im assuming you want to know how many different color arrangements for selecting n balls, n=1,2,3... n=1, 3 ways --> W or G or B n=2 , 6 ways --> 2W or 2G or 2B or WG or WB or GB n=3, 10 ways -->3W or 3B or 3G or 2W1G or 2W1B or 2G1W or 2G1B or 2B1W or 2B1G or WGB and so on.. n=4, 15 ways

OpenStudy (anonymous):

The answer shuld be 879

OpenStudy (dumbcow):

oh so you have to add them all up all the way to n=26 well there is symmetry, so n=0 is same as n=26 , n=1 same as n=25 and so on ... as long as n is not greater than the number of possible color balls, then pattern is the num of ways equals (n+2) choose 2 \[\sum_{1}^{7}\left(\begin{matrix}n+2 \\ 2\end{matrix}\right) = 119\] then figure out n=8 to 12 , add to 1198 , double it then add n=13

OpenStudy (anonymous):

@experimentX @satellite73

OpenStudy (anonymous):

@mukushla @siddhantsharan

OpenStudy (anonymous):

11*10*8 - 1= 879 should be the answer. I think. Can you check?

OpenStudy (anonymous):

Yes...u r correct...but hw did u do it...

OpenStudy (anonymous):

We can select: No white ball....1 One white ball....2 Two white balls.....3 . . . . 10 white balls. .....11 => Total 11 ways to choose. Similarly 10 ways for green and 8 ways for black. However you will have to subtract the one case where no balls are chosen. Therefore total= (Ways to choose white)*(ways to choose black)*(ways to choose green) - 1 = 11*10*8 -1